OpenStudy (anonymous):

Suppose a triangle has two sides of the length of 3 and 4 and that the angle between the two sides is pi over 3. What is the length of the third side of the triangle?

OpenStudy (jdoe0001):

have you covered the Law of Cosines yet?

OpenStudy (jdoe0001):

 \(\bf \textit{Law of Cosines} = c^2 = a^2+b^2-(2ab)cos(\measuredangle C)\\ \quad \\ \implies c = \sqrt{a^2+b^2-(2ab)cos(\measuredangle C)}\)

OpenStudy (anonymous):

yeah, im doing the law of cosines now. I just wasn't sure how to set It up. thank you!